# Optimal Distributed Similarity Estimation for Unitary Channels

### v1

@@ Appendix D Similarity Estimation with Independent Classical Shadow
In this section, we consider completing DSEU with independent classical shadow estimation for unitary channels (CSEU) [34]. In Appendix D.1, we describe the algorithm to complete DSEU and analyze the corresponding query complexity given in Theorem 5 of the main text. In Appendix D.2, we summarize several lemmas used in the proof.
D.1 Algorithm & Proof of Theorem 5
@@
We first briefly introduce CSEU. Given a unknown unitary channel $\mathcal{U}$, we can obtain the following classical snapshot by inputting state $|{\psi}\rangle^{\otimes s}$ and performing POVM $\mathcal{M}_{s}$ (defined in Eq. (5)),
@@
where $|{\psi}\rangle$ is sampled from a state $4$-design ensemble, and $|{\phi}\rangle$ is the measurement result of POVM $\mathcal{M}_{s}$ . As shown in [34, Eq. (10)], the expectation of $\tilde{X}$ is Choi operator of unitary $U$, i.e.,
@@
We can complete DSEU via CSEU as follows. First, we independently perform CSEU on each quantum devices, and obtain two set of classical snapshots of $\mathcal{U},\mathcal{V}$, labeled as $\{\tilde{X}\}$ and $\{\tilde{Y}\}$. Then, we can estimate the similarity $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$ with these independent classical snapshots. We summarize this algorithm in Algorithm 3. In the following, we prove that this algorithm can indeed complete DSEU and analyze its query complexity.
@@
    TsTs queries of unknown unitary channels $\mathcal{U}$ and $\mathcal{V}$ acting on $d$-dimension Hilbert space $\mathcal{H}$.
Output: an estimation of $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$.
@@
4  Measure $(\mathcal{U}(|{\psi_{t,A}}\rangle\!\langle{\psi_{t,A}}|))^{\otimes s}$ with the POVM $\mathcal{M}_{s}$ and obtains result $|{\phi_{t,A}}\rangle$.
@@
6  Measure $(\mathcal{V}(|{\psi_{t,B}}\rangle\!\langle{\psi_{t,B}}|))^{\otimes s}$ with the POVM $\mathcal{M}_{s}$ and obtains result $|{\phi_{t,B}}\rangle$.
@@
$\tilde{\gamma}=\frac{1}{T^{2}d^{2}}\sum_{i,j=1}^{T}\operatorname{Tr}\left[\tilde{X}_{i}^{\dagger}\tilde{Y}_{j}\right].$ (133)
@@
First, we prove that $\tilde{\gamma}$ is an unbiased estimator of $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$:
$\mathbb{E}\tilde{\gamma}$ $=\frac{1}{d^{2}}\mathbb{E}\operatorname{Tr}\left[\tilde{X}^{\dagger}\tilde{Y}\right$] (134)
$=\frac{1}{d^{2}}\operatorname{Tr}\left[J_{\mathcal{U}}^{\dagger}J_{\mathcal{V}}\right]$ (135)
$=\frac{1}{d^{2}}\sum_{ijkl}\operatorname{Tr}\left[\left(U|{j}\rangle\!\langle{i}|U^{\dagger}\otimes|{j}\rangle\!\langle{i}|\right)\left(V|{k}\rangle\!\langle{l}|V^{\dagger}\otimes|{k}\rangle\!\langle{l}|\right)\right]$ (136)
$=\frac{1}{d^{2}}\sum_{ij}\operatorname{Tr}\left[U|{j}\rangle\!\langle{i}|U^{\dagger}V|{i}\rangle\!\langle{j}|V^{\dagger}\right]$ (137)
$=\frac{\operatorname{Tr}^{2}\left[U^{\dagger}V\right]}{d^{2}}.$ (138)
Then, we analyze the query complexity of this classical shadow based algorithm. The variance of estimator $\tilde{\gamma}$ is given by
@@
$=\frac{1}{T^{4}d^{4}}\mathbb{E}\left(\sum_{i,j,k,l}^{T}\operatorname{Tr}\left[\tilde{X}_{i}^{\dagger}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}^{\dagger}\tilde{Y}_{l}\right]\right)-\frac{\operatorname{Tr}^{4}[U^{\dagger}V]}{d^{4}}.$ (140)
After expanding the expectation in the above equation, there are four terms as follows.
@@
    $i=k$ and $j=l$, there are $T^{2}$ terms and with Lemma 18, we have
    $\mathbb{E}\left(\sum_{i,j}\operatorname{Tr}^{2}\left[\tilde{X}_{i}^{\dagger}\tilde{Y}_{j}\right]\right)=T^{2}\mathbb{E}\operatorname{Tr}^{2}\left[X^{\dagger}Y\right]=T^{2}\mathcal{O}\left(\frac{d^{4}(d+s)^{2}}{s^{2}}\right).$ (141)
@@
    For $i\neq k$ and $j\neq l$ case, there are $T^{2}(T-1)^{2}$ terms and we have
    $\mathbb{E}\left(\sum_{i\neq k,j\neq l}\operatorname{Tr}\left[\tilde{X}_{i}^{\dagger}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}^{\dagger}\tilde{Y}_{l}\right]\right)=T^{2}(T-1)^{2}\operatorname{Tr}^{4}[U^{\dagger}V].$ (142)
@@
    For $i=k$ and $j\neq l$ case, there are $T^{2}(T-1)$ terms and with Lemma 19 we have
    $\mathbb{E}\left(\sum_{i=k,j\neq l}\operatorname{Tr}\left[\tilde{X}_{i}^{\dagger}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}^{\dagger}\tilde{Y}_{l}\right]\right)=T^{2}(T-1)\mathbb{E}\operatorname{Tr}\left[X^{\dagger}\tilde{Y}_{1}\right]\operatorname{Tr}\left[X^{\dagger}\tilde{Y}_{2}\right]\leq\mathcal{O}\left(T^{3}d^{4}\right).$ (143)
@@
    For $i\neq k$ and $j=l$ case, there are $T^{2}(T-1)$ terms and with Lemma 19 we have
    $\mathbb{E}\left(\sum_{i\neq k,j=l}\operatorname{Tr}\left[\tilde{X}_{i}^{\dagger}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}^{\dagger}\tilde{Y}_{l}\right]\right)=T^{2}(T-1)\mathbb{E}\operatorname{Tr}\left[\tilde{X}_{1}^{\dagger}Y\right]\operatorname{Tr}\left[\tilde{X}_{2}^{\dagger}Y\right]\leq\mathcal{O}\left(T^{3}d^{4}\right).$ (144)
@@
Thus, to achieve $\varepsilon$ additive error, we require the query times satisfy
@@
Therefore, we complete the proof of Theorem 5.
@@ D.2 Technical Lemmas
Suppose that $\phi$ and $\psi$ are random input states and measurement outcome of learning unitary channel $\mathcal{U}$, then we have
@@
$\Delta_{U,1}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{1}{\kappa_{2}}\Pi_{\rm sym}^{(2)}\otimes\Pi_{\rm sym}^{(2)},$ (149)
$\Delta_{U,2}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{s}{\kappa_{3}}(I\otimes U\otimes I\otimes I)\left[I_{1}\otimes\left(\Pi_{\rm sym}^{(3)}\right)_{2,3,4}\right](I\otimes U^{\dagger}\otimes I\otimes I)\left(\Pi_{\rm sym}^{(2)}\otimes I\otimes I\right),$ (150)
$\Delta_{U,3}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{s}{\kappa_{3}}(U\otimes I\otimes I\otimes I)\left[I_{2}\otimes\left(\Pi_{\rm sym}^{(3)}\right)_{1,3,4}\right](U^{\dagger}\otimes I\otimes I\otimes I)\left(\Pi_{\rm sym}^{(2)}\otimes I\otimes I\right),$ (151)
$\Delta_{U,4}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{s(s-1)}{2\kappa_{4}}(U\otimes U\otimes I\otimes I)\Pi_{\rm sym}^{(4)}(U^{\dagger}\otimes U^{\dagger}\otimes I\otimes I).$ (152)
@@
Suppose $X$ and $Y$ are classical snapshot defined in Eq. (130) with POVM $\mathcal{M}_{s}$ for unitary channels $\mathcal{U}$ and $\mathcal{V}$, we have
$\mathbb{E}\operatorname{Tr}^{2}\left[X^{\dagger}Y\right]=\mathcal{O}\left(\frac{d^{4}(d+s)^{2}}{s^{2}}\right).$ (153)
@@
With Lemma 10 and the definition of $X$, $Y$, we have
$\mathbb{E}\operatorname{Tr}^{2}\left[X^{\dagger}Y\right]=\frac{1}{s^{2}}\operatorname{Tr}^{2}\left[\left(d(d+1)(d+s){\phi_{1}}\otimes{\psi_{1}}^{T}-(d+1+s)I\otimes I\right)\left(d(d+1)(d+s){\phi_{2}}\otimes{\psi_{2}}^{T}-(d+1+s)I\otimes I\right)\right]$ (154)
@@
We consider the first two terms respectively as follows.
@@
    For the first term, with Lemma 17, there are $16$ terms,
@@
    We analyze these terms as follows
    $\operatorname{Tr}[\Delta_{U,1}\Delta_{V,1}]$ $=\frac{\operatorname{Tr}^{2}\left[\Pi_{\rm sym}^{(2)}\right]}{\kappa_{2}^{2}}=1,$ (157)
@@
    Therefor, for the first term, we have
@@
    Likewise, we have $\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{1}\otimes\psi^{T}_{1}\right]=\mathcal{O}(1)$.
@@
$\mathbb{E}\operatorname{Tr}^{2}\left[X^{\dagger}Y\right]$ $\leq\mathcal{O}\left(\frac{d^{4}}{s^{2}(d+s)^{2}}+\frac{s^{2}}{(d+s)^{2}}\right)+\frac{d^{4}(d+1+s)^{2}}{s^{2}}=\mathcal{O}\left(\frac{d^{4}(d+s)^{2}}{s^{2}}\right).$ (168)
@@
[…]
$\mathbb{E}\operatorname{Tr}\left[X_{1}^{\dagger}Y\right]\operatorname{Tr}\left[X_{2}^{\dagger}Y\right]$ $\leq\mathcal{O}\left(d^{4}\right).$ (169)
[…]
$\mathbb{E}\operatorname{Tr}\left[X^{\dagger}Y_{1}\right]\operatorname{Tr}\left[X^{\dagger}Y_{2}\right]$ $\leq\mathcal{O}\left(d^{4}\right).$ (170)
@@
$\;\mathbb{E}\operatorname{Tr}\left[X_{1}^{\dagger}Y\right]\operatorname{Tr}\left[X_{2}^{\dagger}Y\right]=\operatorname{Tr}\left[\left(\mathbb{E}X\right)^{\otimes 2}\mathbb{E}Y^{\otimes 2}\right]$ (171)
@@
$\mathbb{E}\operatorname{Tr}\left[X_{1}^{\dagger}Y\right]\operatorname{Tr}\left[X_{2}^{\dagger}Y\right]$ $=\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\operatorname{Tr}\left[J_{U}^{\otimes 2}\mathbb{E}\left(\phi_{2}\otimes\psi_{2}^{T}\right)^{\otimes 2}\right]-\frac{2d(d+s+1)}{s}\operatorname{Tr}^{2}\left[U^{\dagger}V\right]-\frac{d^{2}(d+s+1)^{2}}{s^{2}}.$ (175)
@@
$\sum_{i,j,k,l}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\Delta_{V,1}\right]$ $=\frac{s}{\kappa_{2}}\sum_{i,j,k,l}\operatorname{Tr}^{2}\left[\Pi_{\rm sym}^{(2)}|{jl}\rangle\!\langle{ik}|\right]=\frac{sd(d+1)}{2\cdot\kappa_{2}}=s,$ (178)
@@
$\mathbb{E}\operatorname{Tr}\left[X_{1}^{\dagger}Y\right]\operatorname{Tr}\left[X_{2}^{\dagger}Y\right]$ $\leq\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\mathcal{O}\left(\frac{s^{2}}{(d+s)(d+s+1)}\right)=\mathcal{O}\left(d^{4}\right).$ (182)
Likewise for $\mathbb{E}\operatorname{Tr}[X^{\dagger}Y_{1}]\operatorname{Tr}[X^{\dagger}Y_{2}]$. ∎

### v2

@@ Appendix D DSEC with Independent Classical Shadow
In this section, we consider completing DSEC with independent classical shadow estimation for unitary channels (CSEU) [37]. In Appendix D.1, we describe the algorithm for completing DSEC and analyze the corresponding query complexity given in Proposition 6 of the main text. In Appendix D.2, we provide several lemmas used in the proof.
D.1 Algorithm & Proof of Proposition 6
@@
We first briefly introduce CSEU. Given an unknown unitary channel $\mathcal{U}(\cdot)=U(\cdot)U^{\dagger}$, we can obtain the following classical snapshot by inputting the state $|{\psi}\rangle^{\otimes s}$ and performing the POVM $\mathcal{M}_{s}$ (defined in Eq. (4)),
@@
where $|{\psi}\rangle$ is sampled from a state $4$-design ensemble, and $|{\phi}\rangle$ is the measurement result of the POVM $\mathcal{M}_{s}$. Here $\phi\mathrel{\mathop{\ordinarycolon}}=|{\phi}\rangle\!\langle{\phi}|$ is the shorthand of pure state $|{\phi}\rangle$ and similar notation is used for $\psi$. Let $J_{U}$ be the Choi operator of unitary channel $\mathcal{U}$. As shown in [37, Eq. (10)], the expectation of $\tilde{X}$ is $J_{U}$, i.e.,
@@
We can complete DSEC via CSEU as follows. First, we independently perform CSEU on each quantum device, and obtain two sets of classical snapshots of $\mathcal{U}(\cdot)=U(\cdot)U^{\dagger},\mathcal{V}(\cdot)=V(\cdot)V^{\dagger}$, labeled as $\{\tilde{X}\}$ and $\{\tilde{Y}\}$. Then, we can estimate the similarity $\operatorname{Tr}[J_{U}J_{V}]/d^{2}$ with these independent classical snapshots, where $J_{U}$ and $J_{V}$ are the Choi operators of $\mathcal{U}$ and $\mathcal{V}$, respectively. We summarize this algorithm in Algorithm 3. In the following, we prove that this algorithm can indeed complete DSEC and analyze its query complexity.
@@
    TsTs queries to unknown unitary channels $\mathcal{U}$ and $\mathcal{V}$ acting on a $d$-dimensional Hilbert space $\mathcal{H}$.
Output: an estimation of $\operatorname{Tr}[J_{U}J_{V}]/d^{2}$.
@@
4  Measure $(\mathcal{U}(|{\psi_{t,A}}\rangle\!\langle{\psi_{t,A}}|))^{\otimes s}$ with the POVM $\mathcal{M}_{s}$ and obtain the result $|{\phi_{t,A}}\rangle$.
@@
6  Measure $(\mathcal{V}(|{\psi_{t,B}}\rangle\!\langle{\psi_{t,B}}|))^{\otimes s}$ with the POVM $\mathcal{M}_{s}$ and obtain the result $|{\phi_{t,B}}\rangle$.
@@
$\tilde{\gamma}=\frac{1}{T^{2}d^{2}}\sum_{i,j=1}^{T}\operatorname{Tr}\left[\tilde{X}_{i}\tilde{Y}_{j}\right].$ (154)
@@
First, we prove that $\tilde{\gamma}$ is an unbiased estimator of $\operatorname{Tr}[J_{U}J_{V}]/d^{2}$:
$\mathbb{E}\tilde{\gamma}$ $=\frac{1}{d^{2}}\mathbb{E}\operatorname{Tr}\left[\tilde{X}\tilde{Y}\right]=\frac{1}{d^{2}}\operatorname{Tr}\left[J_{U}J_{V}\right].$ (155)
Then, we analyze the query complexity of this classical shadow-based algorithm. The variance of estimator $\tilde{\gamma}$ is given by
@@
$=\frac{1}{T^{4}d^{4}}\mathbb{E}\left(\sum_{i,j,k,l=1}^{T}\operatorname{Tr}\left[\tilde{X}_{i}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}\tilde{Y}_{l}\right]\right)-\frac{\operatorname{Tr}^{4}[U^{\dagger}V]}{d^{4}}.$ (157)
After expanding the expectation in the above equation, there are four terms:
@@
    For $i=k$ and $j=l$, there are $T^{2}$ terms, and using Lemma 20, we have
    $\mathbb{E}\left(\sum_{i,j}\operatorname{Tr}^{2}\left[\tilde{X}_{i}\tilde{Y}_{j}\right]\right)=T^{2}\mathbb{E}\operatorname{Tr}^{2}\left[XY\right]=T^{2}\mathcal{O}\left(\frac{d^{4}(d+s)^{2}}{s^{2}}\right).$ (158)
@@
    For $i\neq k$ and $j\neq l$, there are $T^{2}(T-1)^{2}$ terms, and we have
    $\mathbb{E}\left(\sum_{i\neq k,j\neq l}\operatorname{Tr}\left[\tilde{X}_{i}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}\tilde{Y}_{l}\right]\right)=T^{2}(T-1)^{2}\operatorname{Tr}^{4}[U^{\dagger}V].$ (159)
@@
    For $i=k$ and $j\neq l$, there are $T^{2}(T-1)$ terms, and using Lemma 21 we have
    $\mathbb{E}\left(\sum_{i=k,j\neq l}\operatorname{Tr}\left[\tilde{X}_{i}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}\tilde{Y}_{l}\right]\right)=T^{2}(T-1)\mathbb{E}\operatorname{Tr}\left[X\tilde{Y}_{1}\right]\operatorname{Tr}\left[X\tilde{Y}_{2}\right]\leq\mathcal{O}\left(T^{3}d^{4}\right).$ (160)
@@
    For $i\neq k$ and $j=l$, there are $T^{2}(T-1)$ terms, and using Lemma 21 we have
    $\mathbb{E}\left(\sum_{i\neq k,j=l}\operatorname{Tr}\left[\tilde{X}_{i}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}\tilde{Y}_{l}\right]\right)=T^{2}(T-1)\mathbb{E}\operatorname{Tr}\left[\tilde{X}_{1}Y\right]\operatorname{Tr}\left[\tilde{X}_{2}Y\right]\leq\mathcal{O}\left(T^{3}d^{4}\right).$ (161)
@@
Thus, to achieve $\varepsilon$ additive error, we require the number of queries to satisfy
@@
Therefore, we complete the proof of Proposition 6.
@@ D.2 Technical Lemmas
Suppose that $\phi$ and $\psi$ are the random input states and measurement outcomes of learning a unitary channel $\mathcal{U}$, then we have
@@
$\Delta_{U,1}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{1}{\kappa_{2}}\Pi_{\rm sym}^{(d,2)}\otimes\Pi_{\rm sym}^{(d,2)},$ (166)
$\Delta_{U,2}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{s}{\kappa_{3}}(I\otimes U\otimes I\otimes I)\left[I_{1}\otimes\left(\Pi_{\rm sym}^{(d,3)}\right)_{2,3,4}\right](I\otimes U^{\dagger}\otimes I\otimes I)\left(\Pi_{\rm sym}^{(2)}\otimes I\otimes I\right),$ (167)
$\Delta_{U,3}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{s}{\kappa_{3}}(U\otimes I\otimes I\otimes I)\left[I_{2}\otimes\left(\Pi_{\rm sym}^{(d,3)}\right)_{1,3,4}\right](U^{\dagger}\otimes I\otimes I\otimes I)\left(\Pi_{\rm sym}^{(d,2)}\otimes I\otimes I\right),$ (168)
$\Delta_{U,4}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{s(s-1)}{2\kappa_{4}}(U\otimes U\otimes I\otimes I)\Pi_{\rm sym}^{(d,4)}(U^{\dagger}\otimes U^{\dagger}\otimes I\otimes I).$ (169)
@@
Let $X$ and $Y$ be classical snapshots defined in Eq. (151) with POVM $\mathcal{M}_{s}$ for unitary channels $\mathcal{U}$ and $\mathcal{V}$, we have
$\mathbb{E}\operatorname{Tr}^{2}\left[XY\right]\leq\mathcal{O}\left(\frac{d^{4}(d+s)^{2}}{s^{2}}\right).$ (170)
@@
Using Lemma 9 and the definition of $X,Y$, we have
$\mathbb{E}\operatorname{Tr}^{2}\left[XY\right]=\frac{1}{s^{2}}\operatorname{Tr}^{2}\left[\left(d(d+1)(d+s){\phi_{1}}\otimes{\psi_{1}}^{T}-(d+1+s)I\otimes I\right)\left(d(d+1)(d+s){\phi_{2}}\otimes{\psi_{2}}^{T}-(d+1+s)I\otimes I\right)\right]$ (171)
@@
We consider the first two terms respectively as follows:
@@
    For the first term, with Lemma 19, there are 1616 terms:
@@
    We analyze these terms as follows:
    $\operatorname{Tr}[\Delta_{U,1}\Delta_{V,1}]$ $=\frac{\operatorname{Tr}^{2}\left[\Pi_{\rm sym}^{(d,2)}\right]}{\kappa_{2}^{2}}=1,$ (174)
@@
    Therefore, for the first term, we have
@@
    Likewise, $\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{2}\otimes\psi^{T}_{2}\right]=\mathcal{O}(1)$.
@@
$\mathbb{E}\operatorname{Tr}^{2}\left[XY\right]$ $\leq\mathcal{O}\left(\frac{d^{4}}{s^{2}(d+s)^{2}}+\frac{s^{2}}{(d+s)^{2}}\right)+\frac{d^{4}(d+1+s)^{2}}{s^{2}}=\mathcal{O}\left(\frac{d^{4}(d+s)^{2}}{s^{2}}\right).$ (185)
@@
Let $X_{1}$ and $X_{2}$ be classical snapshots defined in Eq. (151) with POVM $\mathcal{M}_{s}$ for unitary channel $\mathcal{U}$, and $Y$ be a classical snapshot for unitary channel $\mathcal{V}$, we have
$\mathbb{E}\operatorname{Tr}\left[X_{1}Y\right]\operatorname{Tr}\left[X_{2}Y\right]$ $\leq\mathcal{O}\left(d^{4}\right).$ (186)
Let $Y_{1}$ and $Y_{2}$ be classical snapshots defined in Eq. (151) with POVM $\mathcal{M}_{s}$ for unitary channel $\mathcal{V}$, and $X$ be a classical snapshot for unitary channel $\mathcal{U}$, we have
$\mathbb{E}\operatorname{Tr}\left[XY_{1}\right]\operatorname{Tr}\left[XY_{2}\right]$ $\leq\mathcal{O}\left(d^{4}\right).$ (187)
@@
$\;\mathbb{E}\operatorname{Tr}\left[X_{1}Y\right]\operatorname{Tr}\left[X_{2}Y\right]=\operatorname{Tr}\left[\left(\mathbb{E}X\right)^{\otimes 2}\mathbb{E}Y^{\otimes 2}\right]$ (188)
@@
$\mathbb{E}\operatorname{Tr}\left[X_{1}Y\right]\operatorname{Tr}\left[X_{2}Y\right]$ $=\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\operatorname{Tr}\left[J_{U}^{\otimes 2}\mathbb{E}\left(\phi_{2}\otimes\psi_{2}^{T}\right)^{\otimes 2}\right]-\frac{2d(d+s+1)}{s}\operatorname{Tr}^{2}\left[U^{\dagger}V\right]-\frac{d^{2}(d+s+1)^{2}}{s^{2}}.$ (192)
@@
$\sum_{i,j,k,l}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\Delta_{V,1}\right]$ $=\frac{s}{\kappa_{2}}\sum_{i,j,k,l}\operatorname{Tr}^{2}\left[\Pi_{\rm sym}^{(d,2)}|{jl}\rangle\!\langle{ik}|\right]=\frac{sd(d+1)}{2\cdot\kappa_{2}}=s,$ (195)
@@
$\mathbb{E}\operatorname{Tr}\left[X_{1}Y\right]\operatorname{Tr}\left[X_{2}Y\right]$ $\leq\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\mathcal{O}\left(\frac{s^{2}}{(d+s)(d+s+1)}\right)=\mathcal{O}\left(d^{4}\right).$ (199)
The same bound holds for $\mathbb{E}\operatorname{Tr}[XY_{1}]\operatorname{Tr}[XY_{2}]$ by symmetry. ∎